OpenStudy (anonymous):

What is the height of a triangle that has an area of 60 yd2 and a base with a length of 12 yd? A. 0.1 yd B. 2.5 yd C. 5 yd D. 10 yd

OpenStudy (anonymous):

all i know is that its not c lol

OpenStudy (anonymous):

i think its B can someone make sure in right

OpenStudy (anonymous):

\[Area = \frac{ b*h }{ 2 }\] \[\frac{ 2*area }{ b } = h\]

OpenStudy (anonymous):

2 * 60 / 12 = 10yad

OpenStudy (anonymous):

ok now i get it thank you